# Contrastive Preference Optimization: Pushing the Boundaries of LLM Performance in Machine Translation

### v3

@@ Abstract
Moderate-sized large language models (LLMs) – those with 7B or 13B parameters – exhibit promising machine translation (MT) performance. However, they do not match the performance of state-of-the-art conventional encoder-decoder translation models or larger-scale LLMs such as GPT-4 (OpenAI, 2023). In this study, we bridge this performance gap. We first assess the shortcomings of supervised fine-tuning for LLMs in the MT task, emphasizing the quality issues present in the reference data, despite being human-generated. Then, in contrast to supervised fine-tuning which mimics reference translations, we introduce Contrastive Preference Optimization (CPO), a novel approach that trains models to avoid generating adequate but not perfect translations. Applying CPO to ALMA (Xu et al., 2023) models with only 22K parallel sentences and 0.1% parameters yields significant improvements. The resulting model, called ALMA-R, can match or exceed the performance of the WMT competition winners and GPT-4 on WMT’21, WMT’22 and WMT’23 test datasets.111We release our code and models at: https://github.com/fe1ixxu/ALMA.
@@ 1 Introduction
Machine translation (MT) predominantly utilizes transformer encoder-decoder architectures (Vaswani et al., 2017), which is evident in prominent models such as NLLB-200 (NLLB TEAM et al., 2022), M2M100 (Fan et al., 2021), BIBERT (Xu et al., 2021), and MT5 (Xue et al., 2021). However, the emergence of decoder-only large language models (LLMs) such as the GPT series (Brown et al., 2020; OpenAI, 2023), Mistral (Jiang et al., 2023), LLaMA (Touvron et al., 2023a, b), Falcon (Almazrouei et al., 2023), inter alia, which have shown remarkable efficacy in various NLP tasks, which attracts the interest of developing machine translation with these decoder-only LLMs. Recent studies (Zhu et al., 2023a; Jiao et al., 2023b; Hendy et al., 2023; Kocmi et al., 2023; Freitag et al., 2023) indicate that larger LLMs such as GPT-3.5 (175B) and GPT-4 exhibit strong translation abilities. However, the performance of smaller-sized LLMs (7B or 13B) still falls short when compared to conventional translation models (Zhu et al., 2023a).
@@
CPO aims to mitigate two fundamental shortcomings of SFT. First, SFT’s methodology of minimizing the discrepancy between predicted outputs and gold-standard references inherently caps model performance at the quality level of the training data. This limitation is significant, as even human-written data, traditionally considered high-quality, is not immune to quality issues (more details in Section 2). For instance, one may notice that some strong translation models are capable of producing translations superior to the gold reference, as illustrated in Figure 1. Secondly, SFT lacks a mechanism to prevent the model from rejecting mistakes in translations. While strong translation models can produce high-quality translations, they occasionally exhibit minor errors, such as omitting parts of the translation. Preventing the production of these near-perfect but ultimately flawed translation is essential. To overcome these issues, we introduce Contrastive Preference Optimization (CPO) to train the ALMA model using specially curated preference data. After CPO training, the ALMA-R model shows marked improvements, achieving performance levels that match or even surpass those of GPT-4 and WMT competition winners.
@@
Pushing the Performance Boundary of SFT We introduce Contrastive Preference Optimization, which offers advantages in terms of memory efficiency, speed, and, crucially, enhanced effectiveness in improving translation quality. CPO breaks the performance bottleneck inherent in SFT’s reference-mimicking learning process and push the performance boundary of models that have reached saturation through SFT training.
Preference Data We build and release a high-quality preference data for the machine translation area.
@@ 2 Gold or Gilded? Scrutinizing Gold Reference Quality
Models We scrutinize the translation outputs from ALMA-13B-LoRA222ALMA-13B-LoRA is the best 13B translation model in the ALMA families. It initially undergoes full-weight fine-tuning on monolingual data, followed by fine-tuning on high-quality human-written parallel data using low-rank adaptation (LoRA) (Hu et al., 2022)., as well as zero-shot translations from the most recent GPT-4 (gpt-4-1106-preview). To assess the quality of these outputs, we employ two of the latest and largest reference-free models, each with a 10B parameter size and demonstrating very high correlation with human judgements (Freitag et al., 2023). These models are Unbabel/wmt23-cometkiwi-da-xxl (henceforth referred to as KIWI-XXL) (Rei et al., 2023) and Unbabel/XCOMET-XXL (subsequently referred to as XCOMET) (Guerreiro et al., 2023).
@@
Data we consider the high-quality and human-written FLORES-200 dataset (NLLB TEAM et al., 2022), comprising both development and test data, amounting to a total of 2009 samples for each language direction, to compare the gold references with the outputs generated by the models. We employed ALMA-13B-LoRA and GPT-4 to perform translations across five English-centric language pairs, covering both translations from and to English. These pairs include German (de), Czech (cs), Icelandic (is), Chinese (zh), and Russian (ru), with Icelandic (is) categorized as a low-resource language and the others as high-resource languages.
Prompt The prompt employed for generating translations with ALMA models is consistent with the one used in Xu et al. (2023). For GPT-4 translation generation, we follow the guidelines suggested by Hendy et al. (2023). The specifics of these prompts are detailed in Appendix B.
@@
Model Outputs Can Be Better References In Table 1, we present the evaluation scores of KIWI-XXL and XCOMET for the gold references, ALMA-13B-LoRA outputs, and GPT-4 outputs. Additionally, we report Win Ratio, reflecting the proportion of instances where model outputs surpass the gold standard references. These metrics are calculated as an average across five languages. Remarkably, even comparing with the high-quality Flores-200 dataset, the average performance of translation models in xxen translations significantly exceeds that of the references, showing approximately 3-4 point increases in KIWI-XXL and 4-6 point gains in XCOMET. Notably, a significant proportion of outputs are rated higher than the references by KIWI-XXL (e.g., 73.24% for ALMA), with a slightly reduced yet still substantial percentage when assessed using XCOMET (60.17% for ALMA). In the enxx direction, while the overall performance between the translations from reference and two systems is comparable, approximately 40% are still deemed superior to the reference translations.
@@
Motivation: Help The Model Learn Rejection The aforementioned findings illustrate that translations produced by advanced models can sometimes surpass the quality of gold standard references. This raises the question of how to effectively utilize such data. A straightforward approach would involve fine-tuning the model using the source and the superior translations as references. While this could enhance the model’s translation abilities, it does not equip the model with the discernment to identify and avoid generating suboptimal translations, exemplified by the “good but not perfect” translations depicted in Figure 2. Consequently, this situation motivates us to develop a new training objective, which aims to instruct the model in prioritizing the generation of higher-quality translations and rejecting lesser ones, in a style of contrastive learning with hard negative examples (Oord et al., 2018; Chen et al., 2020; He et al., 2020; Robinson et al., 2021; Tan et al., 2023). This objective moves beyond the traditional focus on merely minimizing cross-entropy loss towards the reference.
@@ 3.1 Triplet Preference Data
We here details our methodology for constructing preference data . This dataset is developed using the FLORES-200 data (both development and test sets) and encompasses the same language pairs as discussed in Section 2. For each language pair, the dataset comprises 2009 parallel sentences.
@@ 3.2 Deriving the CPO Objective
where is a pre-trained language (translation) model , is the Sigmoid function, and is a hyperparameter. DPO training can be conducted in a supervised fine-tuning style, as it relies exclusively on labeled preference data and does not require interaction between agents and their environment.
@@
4 Experiments
4.1 Data
Following Section 2, we consider 10 translation directions in the paper: csen, deen, isen, zhen, ruen. Building on the ALMA models’ (Xu et al., 2023) insights that a small quantity of high-quality data can yield impressive translation results, our training dataset is even more compact. As detailed in Section 3.1, our preference training data is derived from the FLORES-200 dataset, a subset of which has been also employed in the training of ALMA models. This results in a total of paired sentences. In addition to preference data assessed by large evaluation models, our dataset incorporates 1K internal human-labeled preference data, containing preferred and dis-preferred translations along with human preference. However, the human-labeled data is limited to just two translation directions: enzh and ende. The details regarding the composition and influence of human-labeled data are explored in Appendix D.444TL;DR: A brief overview of the impact of this human-labeled data suggests a minimal effect. In alignment with Xu et al. (2023), our primary focus is on the test set drawn from WMT’21 for is and WMT’22 for other languages. Additionally, we conduct auxiliary experiments evaluating models on WMT’23, covering six directions: deen, zhen, and ruen.
@@ 4.3 Baselines
SoTA Models In this category, our benchmarks are established against, to the best of our knowledge, the strongest publicly available translation models. We first compare with ALMA-13B-LoRA, recognized as one of the top moderate-size language-model based translation systems, surpassing notable conventional models such as NLLB-54B in both WMT’21 and WMT’22. We also compare our results with TowerInstruct555https://huggingface.co/datasets/Unbabel/TowerBlocks-v0.1., a recently released LLM-based translation model and a contemporary work in the field.666 Note that TowerInstruct has used WMT’22 test data for training, so we exclude it from comparison on the WMT’22 test dataset. Additionally, we evaluate against the zero-shot performance of the latest GPT-4 (gpt-4-1106-preview), currently shown to be the best translation model among all LLM-based translation systems (Xu et al., 2023; Zhang et al., 2023; Zeng et al., 2023; Jiao et al., 2023a). Lastly, we include comparisons with the WMT competition winners, representing the highest standard of translation models within the competition, though it is noted that the winning models vary across different language directions.777The WMT winner systems used for comparison in each direction are provided in Appendix E.
SFT and DPO We also compare different training objectives. Given that CPO is designed to steer learning towards preferred data, a straightforward benchmark is to compare its performance against directly SFT on the same preferred data set. Furthermore, considering that CPO is an evolution of DPO, we also include a comparative analysis with DPO.
@@ 4.4 WMT’21 and WMT’22 Results
We present the primary results for enxx and xxen in Table 2 and Table 3, respectively. Our emphasis is primarily on reference-free evaluation models, due to our analysis in Section 2, which questions the reliability of gold references and highlights that evaluations can be compromised by poor-quality references (Kocmi et al., 2023; Freitag et al., 2023). These models include KIWI-XXL, XCOMET, and a smaller yet popular model, Unbabel/wmt22-cometkiwi-da (hereinafter referred to as KIWI-22). Scores highlighted in bold represent the highest achieved across all systems. For a comprehensive comparison, we also include reference-based evaluations using sacreBLEU (Post, 2018) and COMET-22 (Unbabel/wmt22-comet-da) (Rei et al., 2022) in Appendix A.
Comparing With SoTA Models While ALMA-13B-LoRA ranks as one of the top moderate-size LLM translation models, it slightly trails behind GPT-4 and the WMT competition winners. However, the incorporation of CPO significantly enhances ALMA’s capabilities, bringing its performance to a level that is comparable to or even surpasses that of GPT-4 and WMT winners. For example, ALMA-13B-R achieves an average score of 85.74 on KIWI-XXL and 94.05 on XCOMET for enxx translations. These scores outperform GPT-4, which scores 83.83 on KIWI-XXL and 93.23 on XCOMET, as well as the WMT winners, who score 84.81 on KIWI-XXL and 93.78 on XCOMET.
Comparing With SFT and DPO All training objectives in our study are fine-tuned using the ALMA-13B-LoRA model as a base. In Table 2 and 3, we observe that SFT on preferred data marginally enhances the ALMA model’s translation capability for xxen, and results in a slight deterioration for enxx. Similarly, DPO slightly decreases model performance. In contrast, CPO demonstrates significant improvements across all translation directions.
@@ 5.1 Are Translations Really Better or Just Metric-Preferred?
In our study, since the preferred data is selected by reference-free models and the same models are used for evaluation, we investigate the potential for ’cheating’ in the scoring process. Specifically, we question whether the improved translation scores reflect genuinely better translations or if they simply align more closely with the evaluation model’s preferences. This inquiry is addressed in two parts:
@@
At the method level, we question whether training on metric-preferred data always leads to better scores on that metric, regardless of the method we use. However, the connection is not straightforward; for instance, SFT on preferred data paradoxically results in diminished performance across all three metrics as shown in Table 2.
Consequently, our analysis supports the robustness and validity of using reference-free models like KIWI-XXL and XCOMET both for constructing preference data and for evaluation purposes, underscoring the absence of bias in this approach. Furthermore, Table 5 demonstrates that the choice between using KIWI-XXL, XCOMET, or an ensemble of both has a minimal impact on the results.

### v4

@@ Abstract
Moderate-sized large language models (LLMs) – those with 7B or 13B parameters – exhibit promising machine translation (MT) performance. However, they do not match the performance of state-of-the-art conventional encoder-decoder translation models or larger-scale LLMs such as GPT-4 (OpenAI, 2023). In this study, we bridge this performance gap. We first assess the shortcomings of supervised fine-tuning for LLMs in the MT task, emphasizing the quality issues present in the reference data, despite being human-generated. Then, in contrast to supervised fine-tuning which mimics reference translations, we introduce Contrastive Preference Optimization (CPO), a novel approach that trains models to avoid generating adequate but not perfect translations. Applying CPO to ALMA (Xu et al., 2023) models with only 22K parallel sentences and tuning only 0.1% parameters yields significant improvements. The resulting model, called ALMA-R, can match or exceed the performance of the WMT competition winners and GPT-4 on WMT’21, WMT’22 and WMT’23 test datasets.
@@ 1 Introduction
Machine translation (MT) predominantly utilizes transformer encoder-decoder architectures (Vaswani et al., 2017), which is evident in prominent models such as NLLB-200 (NLLB TEAM et al., 2022), M2M100 (Fan et al., 2021), BiBERT (Xu et al., 2021), and MT5 (Xue et al., 2021). However, the emergence of decoder-only large language models (LLMs) such as the GPT series (Brown et al., 2020; OpenAI, 2023), Mistral (Jiang et al., 2023), LLaMA series (Touvron et al., 2023a, b), Falcon (Almazrouei et al., 2023), inter alia, which have shown remarkable efficacy in various NLP tasks, which attracts the interest of developing machine translation with these decoder-only LLMs. Recent studies (Zhu et al., 2023a; Jiao et al., 2023b; Hendy et al., 2023; Kocmi et al., 2023; Freitag et al., 2023) indicate that larger LLMs such as GPT-3.5 (175B) and GPT-4 exhibit strong translation abilities. However, the performance of smaller-sized LLMs (7B or 13B) still falls short when compared to conventional translation models (Zhu et al., 2023a).
@@
CPO aims to mitigate two fundamental shortcomings of SFT. First, SFT’s methodology of minimizing the discrepancy between predicted outputs and gold-standard references inherently caps model performance at the quality level of the training data. This limitation is significant, as even human-written data, traditionally considered high-quality, is not immune to quality issues (more details in Section 2). For instance, one may notice that some strong translation models are capable of producing translations superior to the gold reference, as illustrated in Figure 1. Secondly, SFT lacks a mechanism to prevent the model from rejecting mistakes in translations. While strong translation models can produce high-quality translations, they occasionally exhibit minor errors, such as omitting parts of the translation. Preventing the production of these near-perfect but ultimately flawed translations is essential. To overcome these issues, we introduce Contrastive Preference Optimization (CPO) to train the ALMA model using specially curated preference data. After CPO training, the ALMA-R model shows marked improvements, achieving performance levels that match or even surpass those of GPT-4 and WMT competition winners.
@@
Pushing the Performance Boundary of SFT We introduce Contrastive Preference Optimization, which offers advantages in terms of memory efficiency, speed, and, crucially, enhanced effectiveness in improving translation quality. CPO breaks the performance bottleneck inherent in SFT’s reference-mimicking learning process and pushes the performance boundary of models that have reached saturation through SFT training.111We release our code and models at: https://github.com/fe1ixxu/ALMA.
Preference Data We build and release a high-quality preference dataset for the machine translation area.
@@ 2 Gold or Gilded? Scrutinizing Gold Reference Quality
We scrutinize the translation outputs from ALMA-13B-LoRA222ALMA-13B-LoRA is the best 13B translation model in the ALMA families. It initially undergoes full-weight fine-tuning on monolingual data, followed by fine-tuning on high-quality human-written parallel data using low-rank adaptation (LoRA) (Hu et al., 2022)., as well as zero-shot translations from the most recent GPT-4 (gpt-4-1106-preview). To assess the quality of these outputs, we employ two of the latest and largest reference-free models, each with a 10B parameter size and demonstrating very high correlation with human judgements (Freitag et al., 2023). These models are Unbabel/wmt23-cometkiwi-da-xxl (henceforth referred to as KIWI-XXL) (Rei et al., 2023) and Unbabel/XCOMET-XXL (subsequently referred to as XCOMET) (Guerreiro et al., 2023).
Data
@@
we consider the high-quality and human-written FLORES-200 dataset (NLLB TEAM et al., 2022), comprising both development and test data, amounting to a total of 2009 samples for each language direction, to compare the gold references with the outputs generated by the models. We employed ALMA-13B-LoRA and GPT-4 to perform translations across five English-centric language pairs, covering both translations from and to English. These pairs include German (de), Czech (cs), Icelandic (is), Chinese (zh), and Russian (ru), with Icelandic (is) categorized as a low-resource language and the others as high-resource languages.
Prompt
Model Outputs Can Be Better References
@@
In Table 1, we present the evaluation scores of KIWI-XXL and XCOMET for the gold references, ALMA-13B-LoRA outputs, and GPT-4 outputs. Additionally, we report Win Ratio, reflecting the proportion of instances where model outputs surpass the gold standard references. These metrics are calculated as an average across five languages. Remarkably, even comparing with the high-quality Flores-200 dataset, the average performance of translation models in xxen translations significantly exceeds that of the references, showing approximately 3-4 point increases in KIWI-XXL and 4-6 point gains in XCOMET. Notably, a significant proportion of outputs are rated higher than the references by KIWI-XXL (e.g., 73.24% for ALMA), with a slightly reduced yet still substantial percentage when assessed using XCOMET (60.17% for ALMA). In the enxx direction, while the overall performance between the translations from reference and two systems is comparable, approximately 40% are still deemed superior to the reference translations.
Motivation: Help The Model Learn Rejection
@@
The aforementioned findings illustrate that translations produced by advanced models can sometimes surpass the quality of gold standard references. This raises the question of how to effectively utilize such data. A straightforward approach would involve fine-tuning the model using the source and the superior translations as references. While this could enhance the model’s translation abilities, it does not equip the model with the discernment to identify and avoid generating suboptimal translations, exemplified by the “good but not perfect” translations depicted in Figure 2. Consequently, this situation motivates us to develop a new training objective, which aims to instruct the model in prioritizing the generation of higher-quality translations and rejecting lesser ones, in a style of contrastive learning with hard negative examples (Oord et al., 2018; Chen et al., 2020; He et al., 2020; Robinson et al., 2021; Tan et al., 2023). This objective moves beyond the traditional focus on merely minimizing cross-entropy loss towards the reference.
@@ 3.1 Triplet Preference Data
We here detail our methodology for constructing preference data . This dataset is developed using the FLORES-200 data (both development and test sets) and encompasses the same language pairs as discussed in Section 2. For each language pair, the dataset comprises 2009 parallel sentences.
@@ 3.2 Deriving the CPO Objective
where is a pre-trained language (translation) model , is the Sigmoid function, and is a hyperparameter. The DPO loss is derived a reparameterization process of the ground-truth reward and the corresponding optimal policy in the Proximal Policy Optimization (PPO) framework (Schulman et al., 2017). As a result, DPO training can be conducted in a supervised fine-tuning style, as it relies exclusively on labeled preference data and does not require interaction between agents and their environment.
@@
4 Experiments
4.1 Data
Following Section 2, we consider 10 translation directions in the paper: csen, deen, isen, zhen, ruen. Building on the ALMA models’ (Xu et al., 2023) insights that a small quantity of high-quality data can yield impressive translation results, our training dataset is even more compact. As detailed in Section 3.1, our preference training data is derived from the FLORES-200 dataset, a subset of which has been also employed in the training of ALMA models. This results in a total of paired sentences. We detail the provenance distribution for each language pair from ALMA-13B-LoRA, GPT4, and reference as presented in Table 2. In addition to preference data assessed by large evaluation models, our dataset incorporates 1K internal human-labeled preference data, containing preferred and dis-preferred translations along with human preference. However, the human-labeled data is limited to just two translation directions: enzh and ende. The details regarding the composition and influence of human-labeled data are explored in Appendix D.444TL;DR: A brief overview of the impact of this human-labeled data suggests a minimal effect. In alignment with Xu et al. (2023), our primary focus is on the test set drawn from WMT’21 for is and WMT’22 for other languages. Additionally, we conduct auxiliary experiments evaluating models on WMT’23, covering six directions: deen, zhen, and ruen.
| ALMA-13B-LoRA | GPT-4 | Reference | |
| ende | 46% | 37% | 17% |
| encs | 32% | 41% | 27% |
| enis | 36% | 40% | 24% |
| enzh | 45% | 35% | 20% |
| enru | 31% | 44% | 25% |
@@ 4.3 Baselines
SoTA Models
In this category, our benchmarks are established against, to the best of our knowledge, the strongest publicly available translation models. We first compare with ALMA-13B-LoRA, recognized as one of the top moderate-size language-model based translation systems, surpassing notable conventional models such as NLLB-54B in both WMT’21 and WMT’22. We also compare our results with TowerInstruct555https://huggingface.co/datasets/Unbabel/TowerBlocks-v0.1., a recently released LLM-based translation model and a contemporary work in the field.666 Note that TowerInstruct has used WMT’22 test data for training, so we exclude it from comparison on the WMT’22 test dataset. Additionally, we evaluate against the zero-shot performance of the latest GPT-4 (gpt-4-1106-preview), currently shown to be the best translation model among all LLM-based translation systems (Xu et al., 2023; Zhang et al., 2023; Zeng et al., 2023; Jiao et al., 2023a). Lastly, we include comparisons with the WMT competition winners, representing the highest standard of translation models within the competition, though it is noted that the winning models vary across different language directions.777The WMT winner systems used for comparison in each direction are provided in Appendix E.
SFT and DPO
We also compare different training objectives. Given that CPO is designed to steer learning towards preferred data, a straightforward benchmark is to compare its performance against directly SFT on the same preferred data set. Furthermore, considering that CPO is an evolution of DPO, we also include a comparative analysis with DPO.
@@ 4.4 WMT’21 and WMT’22 Results
We present the primary results for enxx and xxen in Table 3 and Table 4, respectively. Our emphasis is primarily on reference-free evaluation models, due to our analysis in Section 2, which questions the reliability of gold references and highlights that evaluations can be compromised by poor-quality references (Kocmi et al., 2023; Freitag et al., 2023). However, we are not rejecting the use of reference-based models for evaluation but cautioning the potential pitfalls of poor-quality references. The reference-free models used for evaluation include KIWI-XXL, XCOMET, and a smaller yet popular model, Unbabel/wmt22-cometkiwi-da (hereinafter referred to as KIWI-22). Scores highlighted in bold represent the highest achieved across all systems. For a comprehensive comparison, we also include reference-based evaluations using sacreBLEU (Post, 2018) and COMET-22 (Unbabel/wmt22-comet-da) (Rei et al., 2022) in Appendix A.
Comparing With SoTA Models
While ALMA-13B-LoRA ranks as one of the top moderate-size LLM translation models, it slightly trails behind GPT-4 and the WMT competition winners. However, the incorporation of CPO significantly enhances ALMA’s capabilities, bringing its performance to a level that is comparable to or even surpasses that of GPT-4 and WMT winners. For example, ALMA-13B-R achieves an average score of 85.74 on KIWI-XXL and 94.05 on XCOMET for enxx translations. These scores outperform GPT-4, which scores 83.83 on KIWI-XXL and 93.23 on XCOMET, as well as the WMT winners, who score 84.81 on KIWI-XXL and 93.78 on XCOMET.
Comparing With SFT and DPO
All training objectives in our study are fine-tuned using the ALMA-13B-LoRA model as a base. In Table 3 and 4, we observe that SFT on preferred data marginally enhances the ALMA model’s translation capability for xxen, and results in a slight deterioration for enxx. Similarly, DPO slightly decreases model performance. In contrast, CPO demonstrates significant improvements across all translation directions.
@@ 5.1 Are Translations Really Better or Just Metric-Preferred?
In our study, since the preferred data is selected by reference-free models and the same models are used for evaluation, we investigate the potential for “cheating” in the scoring process. Specifically, we question whether the translations become genuinely better or they simply align more closely with the evaluation model’s preferences. This inquiry is addressed in two parts:
@@
At the method level, we question whether training on metric-preferred data always leads to better scores on that metric, regardless of the method we use. Intriguingly, we observe that generating translations favored by the metric — without true improvement — is not easy. For example, fine-tuning the model solely using DPO or SFT on metric-preferred data can even paradoxically lower its performance on this metric (in Table 3). This prompts us to question whether the improvements observed with CPO, an alternative objective that approximates DPO, when trained on the same data, are merely a result of metric bias. Our stance is that if both DPO and SFT fail to achieve improvements through metric bias, it stands to reason that CPO would similarly not benefit solely from such bias.
@@
5.2 Human Evaluation
The preceding analysis provides indirect evidence underscoring the absence of bias. Here, we incorporate human evaluation as direct proof.
we focused on the zhen direction, which aligns with the example presented in Section 2. We selected 400 samples from a total of 1875 test sentences, each sample including a Chinese source and two English translations, one from our base model ALMA-13B-LoRA and the other from ALMA-13B-R. Four bilingual (English and Chinese) speakers were enlisted to rate each translation on a scale from 0 to 6, as per the methodology outlined in Kocmi et al. (2022). We provide clarity on the evaluation criteria used for scoring in our study:
0: it signifies that the translation is nonsensical, failing to convey any coherent meaning.
2: it indicates that the translation partially preserves the meaning of the source text, albeit with substantial inaccuracies or omissions.
4: it denotes that the translation largely maintains the source text’s meaning, with only minor issues such as slight grammatical errors.
6: it represents a perfect translation, accurately conveying the full meaning of the source text without any errors.
To ensure impartiality, each annotator was assigned 100 samples to score, with the order of the translations randomized to conceal their origin. In Table 7, we report the mean scores, rank position (with rank 1 indicating better translation and rank 2 indicating worse translation since we only have two translations to compare), and win ratio (note that both of them win if there is a tie) of ALMA and ALMA-R:
| Avg. score | Avg. rank | Avg. win ratio (%) | Ties (%) | |
| ALMA-13B-LoRA | 4.86 | 1.60 | 62.50 | 40.30 |
| ALMA-13B-R | 5.16 | 1.40 | 77.80 | 40.30 |
The human evaluation results clearly demonstrate that ALMA-13-R outperforms the original ALMA-13B-LoRA. Consequently, our analysis supports the robustness and validity of using reference-free models like KIWI-XXL and XCOMET both for constructing preference data and for evaluation purposes.